OpenStudy (anonymous):

In parallelogram DEFG, the difference between two adjacent angles is 32 degrees. Which is the measure of the larger angle? Answer 118 degrees 114 degrees 106 degrees 96 degrees

OpenStudy (anonymous):

106 degrees Solve (180 - x) - x = 32 for x x=74, 180-x=106

OpenStudy (anonymous):

let the opposite angles be x and y. x- y = 32 [hence y =x -32] 2x + 2y = 360 [hence 2x + 2(x -32) = 360 Therefore expanding and simplifying... 4x -64 = 360 4x = 364 x = 91 y = 59

OpenStudy (anonymous):

can one of u help me on my next question

OpenStudy (anonymous):

Sure